OpenStudy (anonymous):

@Miracrown @zepdrix @ganeshie8 can you guys please help me

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@aryandecoolest

OpenStudy (anonymous):

r u good at factoring integers and polynomials

OpenStudy (anonymous):

@aryandecoolest

Miracrown (miracrown):

I'll help you. Can I please see your question?

OpenStudy (anonymous):

Which equation would best help solve the following problem? Adina has a rectangular garden that measures 9 m wide by 13 m long. She wants to increase the area to 192 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden? A. 192 = (9)(13) B. 192 = (9)(13 + x^2) C. 192 = (9 – x)(13 – x) D. 192 = (9 + x)(13 + x)

Miracrown (miracrown):

What do you think should be the answer?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

solve this. since it is increased by equal proportions: and area of rectangle is length * breadth (9+x)(13+x)=192 factorise it ....

Miracrown (miracrown):

Let us say that the increment is just x then, area = 192 = (9+x)*(13+x) hence we end up with a quadratic equation in x, which is

Miracrown (miracrown):

(9+x)*(13+x) -192 =0

Miracrown (miracrown):

Yes, D is the answer :)

OpenStudy (anonymous):

further solving will give you. x=-25 and x=3(accepted) ....Not required as per ur question

OpenStudy (anonymous):

@Miracrown can u help me with more please